OpenStudy (anonymous):

Find the common difference, d, in the following arithmetic sequence. 6, 11, 16, 21, … a] -6 b] 6 c] -5 c] 5

OpenStudy (anonymous):

well,you can simply subtract two of those consecutive numbers...for example: 11 - 6 = 5 so the answer is 5 and to check this you can subtract others: 16 - 11 =5 21 - 16 = 5